OpenStudy (anonymous):

solve for x. sqrt--(4x+1) +1=40

OpenStudy (anonymous):

x=380

OpenStudy (anonymous):

\[is \it \sqrt{4x+1} +1 = 40....?\]

OpenStudy (anonymous):

subtract 1 from the both sides &: \[4x+1=39^{2}\] 4x=39^2 -1 solve it for x, please... OK?

OpenStudy (anonymous):

x=1520/4 x=380

OpenStudy (radar):

\[\sqrt{4x+1}+1=40\] \[\sqrt{4x+1}=40-1=39\] Squaring both sides \[4x+1=1521\] 4x=1520 x=380

OpenStudy (anonymous):

if u put 380 in for x. the whole problem adds up to 40

OpenStudy (anonymous):

thanks guys (:

OpenStudy (anonymous):

that's right! you find the value of x that satisfies your original equation :)